### v1

# Graph Residual based Method for Molecular Property Prediction

Abstract
Property prediction of materials has recently been of high interest in the recent years in the field of material science. Various Physics-based and Machine Learning models have already been developed, that can give good results. However, they are not accurate enough and are inadequate for critical applications. The traditional machine learning models try to predict properties based on the features extracted from the molecules, which are not easily available most of the time. In this paper, a recently developed novel Deep Learning method, the Graph Neural Network (GNN), has been applied, allowing us to predict properties directly only the Graph-based structures of the molecules. SMILES (Simplified Molecular Input Line Entry System) representation of the molecules has been used in the present study as input data format, which has been further converted into a graph database, which constitutes the training data. This article highlights the detailed description of the novel GRU-based methodology to map the inputs that have been used. Emphasis on highlighting both the regressive property as well as the classification-based property of the GNN backbone. A detailed description of the Variational Autoencoder (VAE) and the end-to-end learning method has been given to highlight the multi-class multi-label property prediction of the backbone. The results have been compared with standard benchmark datasets as well as some newly developed datasets. All performance metrics which have been used have been clearly defined as well as their reason for choice.
keywords:
GNN , VAE, SMILES, multi-label multi-class classification , GRU
1 Introduction
Quantitative Structure Property Relationship (QSPR) a traditional method developed as early as 1935 [1], has provided an efficient framework to facilitate computation on chemical properties by breaking them down into their smallest blocks. The mathematical model allows the molecules to be broken down according to their chemical, structural, physical, and biological properties. The method of property prediction consists of two steps which involve breaking down the molecule into property descriptors followed by regression on these molecules. The molecular descriptors generate a n-dimensional data with $d=[d_{1},d_{2},\ldots,d_{n}]^{T}$ which indicates the presence of a particular feature as generated by the molecular descriptor.
The molecular descriptors can be 0D,1D,2D as well as 3D as described in the [2, 3]. The most common type of descriptors is 0D descriptors which is typically used in ML models as used in Asheri et al [4] work. The 0D molecular descriptors are simply atoms and molecular properties that are used to describe the molecule as defined by [5]. However, these descriptors don’t contain any information of the structural or connectivity between atoms. Several works can be found using these 0D methods [6]. 1D descriptors have been mostly used till now, they can be thought of as functional sub-groups or the building blocks in a molecule. Most of the works that have been done till now are using this method, however, since these don’t contain any structural information, their accuracy is not up to par. In more recent times, the use of 2D descriptors has been used, which allows the molecular topology to be represented in its true form and introduces several geometry-based features that can be represented in graph format like cyclicity, symmetry, and connectivity between atoms. Lastly, a more modern higher dimensional descriptors [7] which can represent complex molecular geometries and dynamics has been used in cases of [8, 9, 10] however, exploiting such features is computationally expensive.
QSPR modeling is a challenging process as it requires subject expertise to identify the functional subgroups. Automating the process, it requires a careful selection of algorithms and data to identify the information descriptor that is to be used in the ML model. This inherently poses a significant challenge since such descriptors cannot be generalized for all tasks and can be found as trial and error. Due to all the above reasons, alternate ways have come up to extract molecular embeddings or fingerprints directly from the molecular structure rather than from its molecular descriptors. This is where Graph Neural Networks (GNNs) offer efficient methodology. A molecule can be arranged in a graph-like structure from which the deep learning module will directly extract the features. This kind of learning on non-Euclidean datasets [50] is also known as geometric-learning and has been made more easily possible due to the recent development of Pytorch-geometric [11].
GNNs are recognized as a potent tool for extracting knowledge from graph-structured data, showcasing impressive performance across a variety of domains. These domains include social networks, molecular structures, drug screening, chip design, traffic flows, recommendation systems, and knowledge graphs. In various tasks such as node classification, graph classification, graph regression, and link prediction, GNNs consistently achieve state-of-the-art performance. The concept of GNN was first introduced by the paper [12] where they extended the recursive neural networks to give rise to a new set of networks that is more node-focused. The work was further extended by Scarselli et al [13] which gave rise to an end-to-end network of GNN along with modern backpropagation techniques. In this work, the authors provided a detailed mathematical description of the GNN method. One of the major important properties of Graphs and GNNs are that they are size-agnostic i.e., the sample size can differ from one input to another as well as permutation invariant. This property provides a major advantage to GNNs over CNN-based architectures since the size of the input graph i.e. the number of nodes can easily vary whereas in CNNs we must provide a fixed set of input size.
There are various tasks that can be performed by GNNs such as node level tasks as mentioned in [14, 15], edge level tasks such as [16] and Graph level tasks as described in [17, 18]. The problem of Molecule property prediction which we are trying to solve is a Graph level task where each molecular graph generates an embedding which is to be further used. The most primitive GNN used for solving such type of problem has been in the form of Message Passing Neural Network (MPNN) [19]. The MPNN network is based on the simple Graph Convolution format which allows each node to learn from its subsequent neighbouring nodes and that information is then aggregated. Similar, to the Graph Convolutions that are used in MPNN, several other aggregators have been proposed, most notable being Graph Attention Network (GAT) based on the multi-head attention mechanism [20], the recently developed GraphSAGE (Graph Sample and Aggregated) [21], a framework for inductive representation learning on large graphs which samples and aggregates feature from a node’s neighbourhood to learn its embedding, GraphConv mechanism [22] etc. Graph Isomorphism Network (GIN) by Xu eta al [23] provides flexibility in capturing the local and global graph structure, making it suitable for regression tasks.
This paper emphasizes predicting both the regression and classification properties of molecules. It presents results for new datasets and compares them with well-established benchmarks. The classification section specifically examines organic and inorganic compounds, highlighting our research’s effectiveness in drug discovery and predicting molecular properties. Graph classification holds particular significance in the realm of molecular property prediction, where molecular structures are represented as graphs, with atoms as nodes and chemical bonds as edges. In this context, graph classification tasks play a crucial role in predicting various molecular properties, such as toxicity, solubility, bioactivity, and drug efficacy, as mentioned in [24, 25] among others. By leveraging graph-based representations, researchers can capture intricate structural features and relationships inherent in molecular compounds, enabling more accurate and comprehensive predictions. Tasks within molecular property prediction often include binary classification to distinguish between toxic and non-toxic compounds, multiclass classification to categorize compounds into different therapeutic classes or activity levels, and regression to predict quantitative properties like molecular weight or solubility. These tasks are pivotal in drug discovery, materials science, and environmental health.
Jiying et al [26] presented an approach to fine-tuning Graph Neural Networks by framing the problem within an Optimal Transport framework, capitalizing on graph topology. Through the introduction of a masked Optimal Transport problem and the definition of the Masked Wasserstein distance, their GTOT-Tuning (Graph Induced Optimal Transport) framework facilitates efficient knowledge transfer from pre-trained models while preserving local feature invariances. They leverage graph structure to enhance knowledge transfer and offer flexibility for integration with complementary techniques like contrastive learning [27] . However, this method poses a high computational demand and has limited efficacy on small datasets or those with sparse graph structures, as is usually the case with medicinal and molecular data.
Another enhancement on GNNs is the Graph Multiset Transformer (GMT), introduced by Jinheon et al [28] as a graph multi-head attention-based global pooling layer. It captures interactions and relations between nodes of a graph based on its structural dependencies. This ensures that GMT satisfies both injectiveness and permutation invariance. While leading the model to learn more accurate representations of entire graphs by capturing structural dependencies between nodes, training a transformer model such as GMT requires significant computational resources. The inner workings of GMT also capture complex structural dependencies, which are hard to interpret in a physical manner.
Another important context that has been explored is the context of multi-class multi-label classification. Multilabel classification is widely explored in the Image modality, with extensive use of transformer-like encoder-decoder models, as shown by Shichao et al [29] and Shilong et al [30]. Graphical Classification relies heavily on Transformer based approaches as well, with GNN based models lacking severely in performance.
Various methods exist for multi-class multi-label methods on different modalities of input data such as [31, 32]. In this context, we are trying to explore the classification capabilities of the GNNs. The first paper proposed this was a simple GCN structure [33] which was used for multi-label classification.
With the introduction of “Attention is all you need” paper [34] and recent advances in Transformer architecture, Shermukhamedov et al [35] introduced ElemBert – v1 and ElemBert – v2, which employs a transformer mechanism for binary classification utilizing structural information. Their models, known for their flexibility, demonstrate adaptability across various datasets. Through benchmarking, they have showcased the state-of-the-art performance of their models in diverse material property prediction tasks, encompassing both organic and inorganic compounds. Additionally, Heyrati et al [24] proposed BioAct-Het, leveraging a heterogeneous Siamese neural network to capture the intricate relationship between drugs and bioactivity classes, thereby consolidating them within a unified latent space. While this approach enhances the identification of drug discovery markers, it faces limitations in effectively handling the simultaneous classification of multiple labels. Notably, both approaches require separate models for training each label, resulting in heightened training and inference times, as well as increased resource requirements per data point. Alpersein et al [36] introduce a new method called the All SMILES Variational Autoencoder (VAE), focusing on enhancing the optimization of molecular properties through VAEs that work with both SMILES string and graph-based molecular representations. This approach [36] handles the encoding of a molecule’s multiple SMILES strings through a layered arrangement of recurrent neural networks. It aggregates the hidden states corresponding to each atom across these representations, employing an attention mechanism to consolidate them into a comprehensive, fixed-size latent vector. The model’s ability to decode this representation back into a distinct set of SMILES strings for the original molecule allows it to establish a nearly one-to-one relationship between the molecule and its latent representation. This technique effectively addresses the issue of SMILES strings’ non-uniqueness and sidesteps the computational demands associated with graph convolution operations. The utilization of VAEs is particularly advantageous in this context because they are adept at learning intertwined representations of complex data distributions. Additionally, VAEs incorporate a regularization mechanism by enforcing a prior distribution on the latent space, which aids in mitigating overfitting and enhances the model’s ability to generalize.
In this paper, an attempt has been made to introduce a novel architecture GRU-based Residual ECC network to capture the inherent pattern in the molecular graphs and introduce long-range dependencies. A complete model overview is presented along with its mathematical formulation to explain the underlying architecture. Apart from that, a diagrammatic representation of the model in an end-to-end fashion has been presented for a better understanding of the workflow for both regression and classification tasks. The highlight of the work is the novel GRU-based Residual GNN, along with its explainability and how it relates to the underlying chemistry has been presented. In this work, a comparison has been made against the SOTA model, which is present, and the performance on benchmark datasets. A detailed statistical analysis of the classification datasets as well as the data processing, has also been described in this study.
2 Graphical Representation of Molecule
Morgan’s technique [37] for generating unique machine descriptions and the establishment of the CAS ONLINE search system were significant developments for the description of the chemical structure database. The widespread adoption of computers in the field of chemical nomenclature led to the prevalent use of line notation. Instead of drawing detailed molecular structures, line notation uses lines, symbols, and characters to convey essential structural information. This is attributed to the computational efficiency with which computers can handle linear data strings [38]. Rapid advancements in computer technology have outpaced the incremental growth of chemical information. The synergy between computer capabilities and chemical knowledge allows for the storage of all existing chemical information on current hardware. The historical challenge of merely storing information in machines has been largely overcome. Current and future efforts are now focused on creating highly efficient systems that deliver chemically relevant information when required. To achieve this, a new chemical language and associated computer programs are being developed, grounded in the innovative information system known as SMILES [39].
The SMILES system was created to be genuinely interactive with computers. Its ease of use relies on computer programs that meticulously reformat the input from the chemical user. This achievement stemmed from the pursuit of specific original objectives, including the unique description of a chemical structure graph. This system was designed to overcome difficulties and provide a user-friendly experience by leveraging computer interactivity and stringent input-recoding processes.
SMILES is a line notation used to encode molecular structures, allowing for a linear representation of the molecule so that a computer program can understand the structure. It specifically represents a valence model of a molecule and is widely used in chemistry for representing molecular structures in a compact and human-readable form. The graph of a chemical structure was to be uniquely described, including but not limited to the molecular graph comprising nodes (atoms) and edges (bond) [40]. Since most of the datasets, comes in SMILES [39] encoded format, it needs to be converted into graph database for use in the GNN. A molecular graph can be represented in the form of $G(n)=(V,E)$ where, $n$ is the number of nodes or vertices i.e the atoms in the molecule and $e_{v,w}$ represents the edges i.e the atomic bonds in the molecule where $e_{v,w}\epsilon E$. A typical example of the molecule in this format has been shown in the Figure 1.
Figure 1: Schematic representation of preprocessing of SMILES data.
Table 1: Representation of encoding of node features
Atom Type Atomic Number Neighbour Hydrogen Formal Charge
Node 1 C 6 1 3 0
Node 2 C 6 3 0 0
Node 3 O 8 1 0 0
Node 4 C 6 1 3 0
Table 2: Representation of encoding of edge features
Bond Conjugated Aromatic Is In Ring Connects
Edge 1 Single No No No (1,2)
Edge 2 Double No No No (2,3)
Edge 3 Single No No No (2,4)
Table 3: Node features
Node Features Description/Exemplary Values Number of features Encodings
Atom Type All possible atom list in the periodic table 117 [0-116]
Bonds Number of bonds the atom is involved in 9 [117-125]
Charge Formal Charge of the atom 1 [126]
Hybridization sp, sp2, sp3, sp3d, sp3d2, sp2d, s, other, unspecified 9 [127-135]
Is Aromatic Indicate if the atom forms any aromatic system 1 [136]
Is in Ring Indicate if the atom forms any ring 1 [137]
Atomic Number Atomic Number 1 [138]
Hs Number of bonded hydrogen atoms 9 [139-147]
Table 4: Edge features
Edge Features Description/Exemplary Values Number of Features Encodings
Bond Type Single, Double, Triple and Aromatic 4 [0-3]
Conjugated Whether the bond is conjugated 1 [4]
Is in Ring Whether the bond forms an aromatic ring 1 [5]
There are several types of atomic as well as bond features present, however, for our study, we use only the basic features that will be available without experimentation to allow for the generalizability of the model. Based on Figure 1 a sample encoding structure of the features are presented in the Tables 1 and 2 along with their encodings. These numerical encodings are then converted to one-hot encoded vectors for use in the geometric dataset. The tables show a small sample of the node and edge feature set.
A complete representation of the number of number of features used for both the node and edge cases respectively, has been shown in Tables 3 and 4. They show the size of the feature space of the one-hot encoded vectors. Proper dimension size along with the type of features with detailed descriptions are presented here.
3 Model Description
3.1 GNN Block
A significant amount of work in the field of GNN, [41, 42] focuses on the utilization of node features to find different patterns and relations between different molecule structures, but fails to consider the edge features, which are particularly important when talking about organic molecules, as information such as number of bonds and aromaticity are an important influence on various properties of organic molecules.
Thus, we consider Edge Conditioned Graph Convolution Neural Networks [43], which modifies the normal message passing algorithms of the traditional GNN to include a correction made by using the relevant edge features- aromaticity, multiple bond presence, resonance, etc. Using an Artificial Neural Network, we obtain representations of the edge features in the feature space of the node features, and then add them to the node features during message passing algorithm, ensuring that the resulting node features obtained have information not just from the surrounding neighbouring nodes, but from the edge features connecting those nodes as well.
The Eq.1 describes the convolution operation of the ECC, where $h_{\theta}$ is a multi–layer perceptron which takes in the edge features $(e_{ij})$ and dynamically gives the required edge weight. Here, $x$ represents the node features that are passed to the model, and $\theta$ represents the parameters of the model to be updated.
$x_{i}^{\prime}=\theta x_{i}+\sum_{j\in N(i)}x_{j}\cdot h_{\theta}(e_{ij})$ (1)
While this solves the inherent problems within implementations of our GNN, we optimize the training with several other methods. It has been observed that many molecules have similar substructures such as aldehyde groups, benzene rings, and other common elements. We believe that since these groups often act in a fairly similar manner, contributing to properties such as solubility, boiling point, etc. in a standard manner across different molecules, a mechanism has been introduced to store these influences in a separate state so that this information can travel through various molecules. Bigger molecules will thus contribute information that can act as a template for smaller molecules, resulting in faster computation efficiency, and a more focused learning. GGNN (Gated Graph Neural Network) [18] tries to store temporal states by induction of a GRU unit in place of each graph node of a GNN during the message passing layer. Since we need a spatial log of the states that are observed by the algorithm, we apply a GRU layer each at the end of the “hops”, i.e. the end of the message passing algorithm, saving the state as needed to preserve node properties across different molecules. The GRU update steps at the $l^{th}$ layer are
[…]
Figure 2: Schematic diagram of the flow of our GNN model.
This approach clearly delineates the flow of information from one layer to the next, emphasizing the sequential dependency between the ECC output and GRU input within the same layer, and the GRU output serving as the ECC input for the subsequent layer.
A pictorial representation of our model is shown in Figure 2 which describes in detail the various parts of our model. The figure shows a combined diagram of our GNN model along with the attached VAE model for classification purposes. As evident from Figure 2 and 9 layers of GNN blocks are added which consist of an ECC layer, Batch Normalization layer, and GRU cell. The number of layers added to the model represents the number of k hops distance from the node, the model aggregates information from. In this case, since we added 3 layers, we are considering each node to aggregate information 3 nodes away. This is especially useful in the case of larger molecules.
However, in our case, too we experiment with deeper models. Experimentation has been done with upto 5 hop structure, but it did not lead to better accuracy of our results, rather the model performed worse. Kipf and Welling [44] showed that increasing the number of hops could decrease the performance due to the inclusion of noise from the additional hop structures. To solve this problem similar to Residual connections in Computer Vision, several methods have been proposed e.g. Highway GCN by Rashimi et al [45], and Jump Knowledge Network by Xu et al [46] .
In this implementation, we use a simple Residual network connection, deriving inspiration from the DeepGCN model of Li et al [47], who used Residual Connections from the Resnet model [48]. A simple version of ResGCN [49] is used in this case where we have a single residual block, which aggregates information from the first hop and second hop which is to be fed in the third hop. Since all molecular graphs are not large, deeper layers are not used. However, a proper accurate comparison of Residual graph networks on molecular structure will be included in a future study.
Additionally, we apply a GraphConv Mechanism [22] described in Eq. 3 after the third hop, with pooling at the top of it. This transforms the node and edge features into a feature space where all the edge features are 1, thus filtering all the remaining information of the edge features into the node features. This is then passed through a pooling layer to obtain a representation of the molecule that can be passed through any subsequent ANN for classification or regression tasks. The implementation of the same is described by
$x_{i}^{\prime}=W_{1}x_{i}+W_{2}\sum_{j\in N(i)}e_{ij}x_{j},$ (3)
where $W_{1}$ and $W_{2}$ are the parameters of the parameters and $e_{ji}$ is the edge weight which by default we consider it 1. The $x$ here represents the node features computed from the successive ECC blocks.
The ANN defined is a simple Multilayer Perceptron consisting of 3 hidden layers of 128 dimensions each and an output layer. Based on the Regression task the output layer dimension is 1 and for the classification task, the output layer dimension is 256 since the embeddings generated by the model are passed to the VAE as described in Sec. 3.2.
3.2 Explanation of the GNN algorithm
In the current model, we add a GRU layer after each hop. The addition of the GRU layer along with residual connection is what increases the performance of the model. A logical explanation along with outputs of some explainers like GNNExplainer [50], and PGExplainer [51] has been presented as proofs of the working of the model.
As can be understood from the input, the molecular graphs are dynamic in nature, i.e. their size varies. Due to this dynamic nature of input, the model encounters molecules of different lengths during training and testing. Despite training the model on a random selection of training set, the model may find it difficult to adjust the weights to a varying set of inputs. Thus, to solve this problem and allow better adaptability to the molecules we implement GRU layers after each hop along with residual connection between the first and the third hop.
To understand the detailed working and effectiveness of the GRU and residual connections, GNNExplainer [50] has been used to present an explanation on the behaviour of the model for selected target property. For this we select two datasets, Lipophilicity [52] and Aqueous solubility [52].
Figure 3: Comparative Representation of the outputs of GNNExplainer on the target Molecule Name: 5-ethylsulfanyl-3-methyl-1-(2-methylpropyl)-6-[[2-(trifluoromethyl)phenyl]methyl] thieno [2,3-d]pyrimidine-2,4-dione, (a) Molecular Structure, (b) Node Feature Importance, (c) Output generated for the base ECC model, (d) Output generated by our ECC model. In this case the target property considered is Lipophilicity.
Figure 4: Comparative Representation of the outputs of GNNExplainer on the target Molecule Name: 2-[2-[2,4-dimethoxy-5-[(2-methyl-3,4-dihydro-2H-quinolin-1-yl)sulfonyl]anilino]-2-oxoethyl]sulfanylacetic acid, (a) Molecular Structure, (b) Node Feature Importance, (c) Output generated for the base ECC model, (d) Output generated by our ECC model. In this case the target property considered is Lipophilicity.
Figure 5: Comparative Representation of the outputs of GNNExplainer on the target Molecule Name: 4-methoxy-1H-indazole (COc1cccc2[nH]ncc12), (a) Molecular Structure, (b) Node Feature Importance, (c) Output generated for the base ECC model, (d) Output generated by our ECC model. In this case the target property considered is Lipophilicity.
Figure 6: Comparative Representation of the outputs of GNNExplainer on the target Molecule Name: (4-methylphenyl)urea, (a) Molecular Structure, (b) Node Feature Importance, ,(c) Output generated for the base ECC model, (d) Output generated by our ECC model. In this case the target property considered is Aqueous Solubility
Figure 7: Comparative Representation of the outputs of GNNExplainer on the target Molecule Name: 6-(hydroxymethyl)-4-methoxyoxane-2,3,5-triol, (a) Molecular Structure, (b) Node Feature Importance, (c) Output generated for the base ECC model, (d) Output generated by our ECC model. In this case the target property considered is Aqueous Solubility
Figure 8: Comparative Representation of the outputs of GNNExplainer on the target Molecule Name: 4-methoxy-1H-indazole (a) Molecular Structure, (b) Node Feature Importance, (c) Output generated for the base ECC model, (d) Output generated by our ECC model. In this case the target property considered is Aqueous Solubility.
The Figure 5(c), 5(d) and Figure 6(c), 6(d) show the results of output of the GNNExplainer model. The actual structure of the molecules along with the node importance has been presented in Figure 5(a) and 5(b) which are the top left and top right. The diagrams in 5(c) and 5(d) in the bottom left and bottom right compare the explanation of our model vs the base ECC model as mentioned by [43]. The strength of the node as well as the edge connections are presented by varying the degrees of the black. The atoms as well as the rings are colour-coded to match the molecules.
Lipophilicity and aqueous solubility are properties that describe the degree of solubility of the input substance in non-polar and polar solvents respectively. As the molecules are predominantly organic in nature, the feature importance indicates carbon and nitrogen to be the most important contributing atoms to the determination of these properties. Features such as formal charge and aromaticity are also deemed as important features, as these affect the polarity of the molecule by determining the final charge and polarity of the molecule due to the probabilistic flow of electrons. This, in turn, directly relates to the fact that the more polar a molecule is, the higher the expected solubility in a polar solvent, while high polarity indicates a low solubility in non-polar solvents. An atom (node) being part of a ring is also identified as a feature of importance since that indicates the presence of a cyclic structure which can significantly impact a molecule’s overall shape, electronic distribution, and thus its solubility properties. Cyclic compounds, especially those that are aromatic, tend to have de–localized $\pi$ electrons which contribute to the molecule’s stability and reactivity. This de-localization can affect the molecule’s polarity. Overall, we can determine the interplay of these factors—atom type, formal charge, aromaticity, and structural features such as being part of a ring—determines the lipophilicity and aqueous solubility of organic molecules.
From the explanation graphs of lipophilicity and aqueous solubility, as shown in Figure 5(d) and 8(d), one observation to be noted is that our model focuses intensively on connections that signify a high amount of electron flow. This is visible by the lighter (weaker) connections between bonds of similar atoms, like C-C bonds, and a darker (stronger) connection between bonds of Carbon with other atoms like Oxygen, Nitrogen, Fluorine, etc., which use their inherent electronegativity to draw electrons towards themselves, creating polar bonds. This polarization within the molecule is a key factor in determining its solubility characteristics. Bonds between atoms of different electronegativities (such as C-O, C-N, or C-F bonds) are more polar, making the molecules containing these bonds more likely to interact with polar solvents and less likely to interact with non-polar molecules using dipole-dipole interactions.
In the example of the molecule in Figure 3(a) the performance of our model against the base ECC [43] model, it can be seen that our model accurately captures relations relevant to solubility that are missed by the base ECC models, such as connections within the $CF_{3}$ subgroup, as well as connections within the neighbouring benzene ring. Additionally, our model captures other important polar bonds that affect lipophilicity, which the ECC model fails to capture. Thus, this ablation study provides concrete proof of the validity of our novel ideation. This also explains how having a balance of so-called electron-giving and electron-receiving groups results in a very high lipophilicity. If we observe Figure 6(c) and 6(d), for the molecule (4-methylphenyl) urea we observe that our model captures heterogeneous bonds that are relevant in spreading the electron charge and making the molecule polar, while the base ECC [43] model highlights no such information. Thus, our model accurately captures the reasoning behind the high value of the aqueous solubility of the molecule.
3.3 Variational Autoencoder (VAE)
Numerous domains, including network system monitoring, proteomic studies, and drug molecule analysis, necessitate the classification of multiple labels based on graph-connected features such as molecule structures and network graphs. Datasets encoded in (SMILES) [39] format, such as Tox21 [52] and Sider [52], encapsulate valuable information pertaining to drug toxicities and other molecule parameters. An initial attempt to establish an enhanced benchmark involved a straightforward model consisting of a single layer of sigmoid neurons appended to the end of a novel GNN, its competitiveness was noteworthy. However, for multilabel prediction tasks, the pertinence of VAEs surpasses conventional models, inspiring a closer exploration of their efficacy.
VAEs operate by encoding input data into a probabilistic latent space and decoding it to reconstruct the input. This reconstruction allows us to align the feature space and the label space, allowing for much more accurate predictions in the multi-label classification scenario.
The utilization of a VAE for capturing the probabilistic latent space presents significant advantages. Firstly, it grants enhanced control over the latent space [15]. Notably, within various AE models, the label-encoder-decoder pathway often outperforms the feature-encoder-decoder pathway. Incorporating the VAE structure within the latent space aids in balancing the complexity of learning and aligning these distinct subspaces. Secondly, achieving smoothness within the latent space is a common objective [16]. VAEs, being probabilistic models, inherently induce local-scale smoothness as the decoder interprets samples rather than specific embeddings. Thirdly, VAE models and their variations exhibit an ability to learn representations characterized by disentangled factors [17]. When both latent spaces, pertaining to features and labels, learn these disentangled factors, it not only facilitates the alignment of these spaces but also contributes significantly to the decoding process, augmenting our results.
Figure 9: Architecture of VAE model used.
While VAEs are often used for generative tasks, the Contrastive Gaussian Mixture Variational Autoencoder (C-GMVAE) [53] is specially designed for multi-label classification (MLC). This model addresses the common challenges faced by vanilla VAEs [54] such as over-regularization, posterior collapse, and label correlation in MLC tasks. The C-GMVAE leverages contrastive learning to strengthen label embedding learning and integrates a Gaussian mixture variational autoencoder (GMVAE) to learn a multimodal latent space synergistically.
The C-GMVAE model comprises two primary components: the Gaussian mixture VAE and the contrastive learning module. The Gaussian mixture VAE adheres to the conventional VAE framework, combining the posterior distribution with a parameter-free isotropic Gaussian prior. During training, it optimizes two essential losses: the KL divergence from the prior to the posterior and the distance between the reconstructed targets and the real targets. This component empowers the model to acquire a multimodal latent space, facilitating probabilistic modeling. A representative diagram of the C-GMVAE model has been presented in Figure 9 to highlight its workflow in the multi-class multi-label classification setting.
A pivotal innovation within the C-GMVAE model is the contrastive learning module. This component employs a contrastive loss to enhance label embedding learning, effectively capturing label information in a supervised multi-label classification (MLC) scenario. By defining anchor, positive, and negative samples, the contrastive loss harnesses label information with great efficacy, eliminating the need for resource-intensive label correlation modules such as covariance matrices and GNNs. This innovative approach underscores the adaptability of contrastive learning, conventionally employed in self-supervised learning, to effectively capture label information in supervised MLC tasks.
3.4 End to End Learning
Since there are two problems to be solved, regression and classification, both will have their own independent learning schemes. The entire GNN backbone can be represented as $F_{n}:G(n)\rightarrow m_{f}$, where $F_{n}$ is the GNN function that generates the molecular fingerprint/embeddings $(m_{f})$. The MLP as defined in Sec.3.2 is automatically integrated in the GNN function which allows for backpropagation. The embeddings are then used for the general downstream tasks. For the regression task, we can define the task in the format $MLP(m_{f})\rightarrow\text{output prediction}$.
Figure 10: Diagrammatic representation of end-to-end learning.
The multilayer-perceptron is applied to molecular embeddings generated by the GNN. Since the entire updation takes place in a single setting and is bound by a single computation graph, proper end-to-end learning takes place. For the multi-class multi-label classification setting, a diagrammatic representation of the workflow has been represented in Figure 10. As can be seen, the SMILES data is first converted to a graph database by PyTorch-geometric [11], with the features of the nodes and edges being extracted by Rdkit [55]. The dataset is then passed through the GNN as represented above to generate $m_{f}$.The VAE part of the execution can be represented as $VAE(m_{f})\rightarrow\hat{p}$ and $\sigma(\hat{p})\rightarrow\text{predicted vector}$.
In the current scenario, an MLP is used as the encoder and the decoder of the VAE, however, it can be easily replaced with much more sophisticated VAE methods such as [56]. The VAE generates a probability vector which when passed through the sigmoid function will give us the predicted vector. In both the regression and classification scenarios, all the entities are attached to an individual computation graph. Thus, this allows for an end-to-end manner for training where the molecular fingerprint generated is dependent on all the entities and backpropagation is allowed throughout the computation graph.
4 Model Setup and Training
For the model setup, we use the GNN backbone described in Sec. 3.1 for both the regression and classification examples. All the hyper-parameters of the graph convolutions as well as the fully connected layers remain the same. In the case of regression property prediction, Huber loss has been used, a combination of MSE and MAE, to account for the presence of outliers. The resultant loss generated is converted into an exponential loss function.
In the case of Multi-class/Multi-label Classification datasets, the Cross entropy loss function has been used. The GNN backbone along with a sigmoid layer of one-hot encoded classes has been used. To prevent class imbalance from affecting our results, a self-adaptive weighted loss-function has been implemented controlled by Adam optimizer [57] is implemented. This prevents some classes from becoming over-dominant thus increasing our F1 score.
The GNN backbone, in conjunction with VAE, has been used to establish an end-to-end pipeline for training. The hyperparameters of the GNN backbone remain the same, and those of the VAE have been described in Sec. 3.3. To explain, the efficacy of the GNN-VAE model, we also present comparative results of the GNN backbone with the last layer as a sigmoid activation function.
Adam optimizer as well as Cosine annealing scheduler with warm restarts has been used in both the regression as well as classification cases. The batch size is 64, as it is found to be the optimum one, and we have run the model for a maximum of 400 epochs. The learning rate is initialized with a value of 0.001 and a weight decay of 5e-4. An early stopping criterion has a patience value of 20 epochs. The entire dataset is divided into 85% Train set, 10 % Test set, and 5 % Validation set and this has been maintained across all the datasets.
5 Regression
In this section, we have demonstrated our experiments on the regression. The main metrics used for quantifying the results have been presented in Table 5. A detailed description of the data-preprocessing steps of all formats of the dataset has been explained below. For a visual representation of our results parity plots are presented
5.1 Dataset Description
There are several properties such as boiling point, melting point, solubility, etc which are not readily available for use in the case of organic as well as inorganic compounds. These properties take time and resources to be found and extraction of group-contribution of these molecules is not easy to find. For novel compounds, these values are not readily available. Thus, the efficacy of the GNN model to directly extract the property from its SMILES [40] representation is accurately demonstrated based on some of the example properties.
For regression problems, we have taken the datasets used by Asheri et al [4] which contains over 24,259 molecular compounds along with their distinct properties. In order for the model. Among them we have chosen properties such as Boiling Point [4], Acentric factor [4], LogWs [4], LogP [4], Melting Point [4], LD50[4]. Apart from that we have also show-cased results on standard benchmark datasets such as the LipO (lipophilicity) dataset [52].
5.2 Regression Results
Our proposed GNN method demonstrates substantial advancements over prior benchmarks in the field. Building upon the foundational work of Asheri et al. [4] and Rittig et al [43], where R2 score of 0.85 has been reported, our revised GNN model achieved a notable enhancement in performance in the Boiling point dataset[4] both in training and testing phases. As a benchmark, in the test phase, our model showcased a significant improvement, achieving a 10.1% increase in R2 score on average over the single GNN model [43] and 3.41% increase over their ensembled GNN model [43], thus surpassing the previous state-of-the-art results.
To verify our model performance, we have stated the results of various other datasets of Asheri et al [4], and we have obtained competitive or better metrics on those sets. It is observed that the model performs extremely well on those sets where the number of data points is comparatively larger than others. This is inherently due to the fact the GNN models require a comparatively vast amount of data to properly generalize.
Table 5: Prediction Results from Graph Regression
DATASET DATAPOINTS R2 MAE MSE MAPE RMSE
Boiling point 507 0.934 12.375 375.83 2.72% 19.38
Melting Point 686 0.816 30.915 1895.08 9.09% 43.53
Acentric factor 172 0.843 0.0089 0.0002 1.48% 0.0142
LogWs 247 0.883 0.0891 0.0167 2.25% 0.1294
Lipo 420 0.781 0.0745 0.0100 2.6% 0.1004
LogP 1202 0.910 0.0256 0.0011 6.99% 0.0334
Acid Dissolution 160 0.766 0.1245 0.0355 1.64% 0.1885
LD50 472 0.606 0.0591 0.0056 2.20% 0.0749
(a) Lipophilicity
(b) Boiling Point
(c) Melting Point
(d) LogP
(e) LogWs
(f) Acentric Factor
(g) Acid Dissolution Factor
(h) LD50
Figure 11: Parity plot of various datasets of [4] on their test sets as presented
On comparing the results of our model with benchmark datasets, such as Lipophilicity [52], we can see an RMSE score of 0.1 on average on the test set, which is a massive improvement from [58] and [59] which are currently the top benchmarks in the Lipophilicity dataset [52] and reports a RMSE score of 0.587 along with pretraining on additional dataset. It is to be noted that for modeling we have not used any externally added datasets which could have helped in generalization. The results of this paper not only live up to the standards of the current benchmark models but also show the robust generalization capability of our model across all sorts of datasets.
To further highlight the distribution of points and visualize the results, Parity plots of the test set have been provided for all the datasets in Figure 11 to show the grouping of the results on the test set. The actual experimental values have been plotted in the x-axis whereas the predicted values have been plotted in the y-axis. From the plots, it can be seen that the data points lie mostly within the ±10% error line, even for large datasets like Boiling Point, LogP, and LD50 as shown in Figure 11(b), 11(e), 11(h). Since none of the data points have been excluded, the parity plots show that even for cases that are outliers the model performs well and the data points lie within the acceptable limits. However, we can see that the data points for melting points are not that closely packed. This can be reflected in the R2 score as well as the model’s failure to generalize much of the dataset. Apart from the Melting Point dataset 11(c), we can say that more than 98% of the dataset falls below the 10% error range.
Additionally, our method’s success on the LogW, LogP, and Acentric factor datasets highlights its adaptability and generalizability on different datasets. For a clear comparison across all datasets as well as with current benchmark models, we have added all forms of results such as MAE, MSE, and R2 which can be associated with the dataset. A comparison of the number of the data points for the test set
6 Classification
Graph classification plays a pivotal role in the prediction of behaviours of different molecules and has various applications in the domains of drug discovery [25], the generation of new materials [60], and identifying the intrinsic properties of different compounds [24]. Given the superior performance of the new GNN model in regression scenarios,it becomes imperative to subject it to classification tasks, thereby affirming its robustness in discerning intricate patterns across data of different tasks. We perform multi-label multi-class classification, a domain where past experiments have been sparse, proving the model’s ability to handle diverse and complex tasks.
In this study, a comparative analysis has been done between a conventional methodology involving GNN feature embeddings fed into a multi-neuron sigmoid layer and a novel approach with our GNNs as a backbone to a VAE model, as expounded in Sec. 3.3. The latter method facilitates end-to-end learning of both the GNN and VAE blocks at once. A comprehensive comparison is performed to highlight the results and advantages of the novel approach over the conventional approach.
6.1 Dataset Description
For the classification setting, we have focused on those datasets that have multiple classes or sub-tasks and each data point is associated with one of those labels. With this setting in mind, popular datasets have been chosen so that we can benchmark our models against other state-of-the-art algorithms. A detailed description of the datasets used has been presented. The Tox21[52] dataset includes a set of chemical compounds examined for toxicity across 12 different biological targets. With over 8,000 compounds, it provides valuable information for predicting the toxicity and potential adverse effects of various chemicals. Another dataset such as SIDER [52] provides a helpful resource about drug side effects. It gathers info from different sources like clinical trials and medical reports. It helps check if drugs are safe and understand side effects. There are 27 tasks in the dataset which give useful insights into possible side effects of different drugs. Furthermore, we have tried our classification models on higher-dimensional label space methods like Toxcast [52] which consists of more than 8000 compounds and 617 labels. Similarly, with PCBA dataset [52], generated by high throughput screening for biologically active molecules. Like Toxcast this consists of 128-dimension label space and over 400,000 compounds.
6.2 Dataset Processing
Regrettably, several SMILE-encoded datasets earmarked for multilabel multiclass classification exhibited anomalies such as missing labels and vacant entries. While rows with empty entries could be easily excluded, coping with partially missing labels posed a more intricate challenge. Imputing these labels was deemed suboptimal, as it risked compromising the authentic intricacies of the real-world data. Consequently, the preprocessing unfolded in two distinct phases. Firstly, entries with over 50% missing labels were expunged from the dataset. Despite a reduction in data volume, this step ensured the preservation of a certain qualitative threshold, retaining crucial representative elements. Secondly, addressing the unavoidable necessity, missing data was imputed with a constant value (we opted for 0). This step, however, introduced some data corruption and exacerbated the existing class imbalance. Notably, the SIDER [52] dataset remained pristine, devoid of such defects, and was used directly. In contrast, Tox21[52] and ToxCast [52] datasets suffered more significantly from this issue.
Figure 12: Statistical plot of Class imbalance as well as missing values of the datasets (a) SIDER, (b) Tox21, (c) PCBA, (d) ToxCast.
Below shown in Figure 12 is a statistical analysis of the imbalances in the dataset with respect to the labels. As can be seen, there is a huge mismatch between the 1’s and 0’s of all the datasets. One of the major problems in these datasets is the presence of huge number of NAN/missing values. This, problem although not present with sider is predominant in the other datasets such as Tox21[52], PCBA [52], and especially ToxCast [52]. Since imputing all the values with 0 was not deemed ideal, we have also conducted experiments by giving the NAN values a separate class to give them proper representation, the results of which have been discussed. Thus, our problem formulation has been presented for both binary classes as well as multi-class multi-label classification.
6.3 Results of GNN+sigmoid
Before the GNN + VAE model, ablation studies have been performed with the base GNN model. Table 6 describes the results obtained directly from the GNN backbone with the last layer being the sigmoid layer. As can be seen from the table the results are not satisfactory enough with just a single sigmoid layer. The scores obtained however surpass those obtained from the GCN [33] backbone, which further highlights the model’s effectiveness.
Table 6: Results of GNN+sigmoid method
DATASET Macro-AUCROC Ha Total Accuracy Micro-F1
SIDER 0.6234 0.656 0.056 0.6956
Tox21 0.5871 0.811 0.46 0.6133
As can be seen from the results, the primitive method of directly producing the labels is not able to be accurately represented by the GNN model and the accuracy is very low. Thus, for multi-class multi-label settings different options of accurately representing the labels to the embeddings we have presented the GNN+VAE model. A major advantage of this is that the VAE can be directly attached to the GNN backbone.
6.4 Results of GNN+VAE model
The predominant metric for multilabel multiclass classification is the micro-AUC-ROC score, as it is robust against class imbalance in different labels. Previous studies, specifically [35] and [24], utilized a single-model-per-label approach in their work on the Tox21[52] and Sider [52] datasets. While effective for molecular property prediction and drug discovery, this approach proves impractical in high-dimensional label-space scenarios, demanding substantial time and resources. In contrast, our GNN+VAE model for multilabel multiclass classification predicts all labels simultaneously, significantly reducing the resources required for both training and inference. Notably, outperforming the state-of-the-art results reported by [24] on the sider dataset, our model achieves an AUC-ROC score of 93.26%, surpassing their reported 91% average. There is also a discrepancy to be noted. The Bio-Hect [24] model reports metrics on a subset of labels, while our model provides predictions for all labels in the dataset. On the tox21 dataset, our model achieves an AUC-ROC score of 84%, somewhat lower than [35]’s mean score of 96%, attributed to the smaller number of labels in the Tox21 dataset contributing to better learning of single-model-per-label systems. Despite this, our approach proves to be a competitive and resource-efficient alternative.
On distinct 85-5-10 train-validation-test splits, our model demonstrates robust performance on sider, achieving a 76.7% hamming accuracy. Notably, the model excels on Tox21 with an impressive 60% accuracy and a consistent 92% hamming accuracy. The observed discrepancy is due to the lower label dimension in Tox21, contributing to the model’s optimal performance in this setting.
On the PCBA dataset, our model underperforms with an AUC score of 75%, quite lower than the benchmark set by [61] at 87%. In the expansive ToxCast dataset featuring an impressive 617 task labels, our model showcases competitive performance, achieving a notable 78.2% AUC-ROC score. This result narrowly surpasses the prior benchmark established by [62]. A similar approach, the Graph Multiset pooling model for graph classification [28], which, like ours, predicts all labels simultaneously with a single model, reports a considerably lower ROC-AUC score of 65.44%.
The versatility and adaptability of our model are evident in its consistent performance across these diverse datasets. The substantial enhancements in Micro-AUC-ROC scores, particularly in Sider, and the model’s ability to provide competitive results even for a high dimensional label space like that of ToxCast, underscore not only the efficacy of our approach but also its potential for accurate and comprehensive multilabel multiclass classification tasks in various domains. This in turn affirms the robustness and generalizability of the new GNN+VAE model. While our model excels at some tasks, some other datasets have better methods to deal with them individually. The upside to this is that our model performs consistently and competitively on all SMILE-encoded classification datasets with SOTA or near-SOTA performance.
Table 7: Predictions for Multi-label Binary-class Classification
DATASET Macro-AUCROC Ha Total Accuracy Micro-F1 Mico-AUCROC
SIDER 0.8321 0.7812 0.0933 0.8560 0.9326
Tox21 0.7476 0.92 0.31 0.7522 0.8478
ToxCast 0.587 0.67 0.04 0.828 0.7821
PCBA 0.6741 0.735 0.035 0.787 0.7521
MicroAUC is favoured over MacroAUC in multilabel classification due to its ability to provide a balanced evaluation of model performance across all labels. Unlike MacroAUC, MicroAUC considers the contributions of each label equally, making it robust to class imbalance and sensitive to rare classes. By aggregating true positive and false positive rates, MicroAUC offers a comprehensive measure of the model’s discriminative ability, making it well-suited for assessing overall performance in multilabel classification tasks.
Apart from the AUCROC score, other major criteria include accuracy. Though the AUCROC is an important metric, it highlights only the model’s discriminatory power, it provides no information about the accuracy of the prediction.
In a multi-label setting, where instances can belong to multiple classes simultaneously, hamming accuracy $(H_{a})$ as described in Eq. 4 emerges as a more pertinent evaluation metric compared to traditional accuracy. Unlike accuracy, which evaluates the correctness of predictions on all labels collectively, Ha measures the fraction of correctly predicted labels per instance. $H_{a}$ accounts for the inherent complexity of multilabel classification tasks by assessing the model’s ability to correctly predict each label independently, offering a more comprehensive assessment of its predictive capabilities. $H_{a}$ is given by
$H_{a}=\frac{1}{n}\sum_{i=1}^{n}\frac{1}{m}\sum_{j=1}^{m}\delta(y_{ij}=\hat{y}_{ij}),$ (4)
where, $n$ is the number of samples, $m$ is the number of labels, $y_{ij}$ is the true value of the jth label for the ith sample, is the Predicted value of the jth label for the ith sample
A representation of the Hamming Accuracy of the model on the test set has been plotted vs the epochs. This shows the model performance with epochs over the dataset. In Figure 13 both total accuracy scores and Hamming accuracy scores have been presented. The reason for the importance of Ha being more than the total accuracy is due to their formulation as described in Eq. 4. The total accuracy is stricter in the sense that it penalizes the entire set of labels of a data point in case of a mismatch in multi-label settings. Whereas, the HA considers each individual label in a label set as a single entity to be compared.
Figure 13: Plot of Hamming Accuracy vs Epochs to observe the improvement in accuracy of the model
Table 8: Predictions for Multi-label Multi-class Classification
DATASET Macro-AUCROC Ha Total Accuracy Mico-AUCROC
Tox21 0.812 0.8806 0.64 0.8478
ToxCast 0.7491 0.8069 0.063 0.8011
PCBA 0.7577 0.735 0.203 0.7967
Apart from the normal prediction, we have had the idea of setting up separate labels for the missing values instead of directly imputing the values. This is highly effective for those datasets where the number of missing values is quite high in proportion to the other classes such as PCBA and ToxCast as can be seen from the statistical analysis in Figure 12 in Sec.6.2. The problem formulation now becomes a 3-class problem consisting of labels 0,1,2 which are then one-hot encoded in the vector space. The reason for adding a separate label is because some of these labels being missing may be inherently correlated with the property of the molecule, i.e. maybe those tests cannot be performed on that molecule to find the property.
From the results, as shown in Table 8, a major increase in R2 values can be seen for Tox21 as well as ToxCast datasets. Furthermore, a substantial increase in total accuracy as well as Hamming accuracy can be seen. Since nothing specific has been provided on the kind of preprocessing that has been done in the papers [26], from the results thus it can be concluded that adding a separate class for the missing values seems more appropriate than imputing them.
7 Conclusions
This work presents a new GRU-based Residual Graph model to overcome the limitations of the base ECC model. An end-to-end learning of the model has been presented on various SMILE-encoded molecular datasets, for both classification and regression tasks. From the Comparison of our model against the current SOTA approaches it has been observed that the model provides adequate improvement from the current SOTA, especially in cases of Lipophilicity and Boiling Point datasets. Apart from that, special focus has been placed in the context of multi-class multi-label classification of various datasets like SIDER, ToxCast and Tox21. Statistical analysis of the classification datasets including missing values and class imbalances have been presented. Both the methods of imputing missing values with 0 and adding them as a separate class have been tried. A quick comparison of the results shows that adding separate labels to the missing values helps improve the classification metrics. A VAE with the novel GNN as a backbone has been presented in this context, which has increased the accuracy of the association of labels with molecular feature embeddings. Our model performs quite competitively in this setting as well. The interpretability of our model and its relationship with the underlying chemistry has been shown here by the outputs of the GNNExplainer algorithm. Our model adheres better to the notion of the underlying chemistry than the base ECC model, as explained by the output graph due to the residual connections and GRU layers. A comprehensive study on the impact of residual connections, the number of layers, and its impact on property prediction will be undertaken as a future work.
Data availability
All the datasets that have been used have been mentioned and are publicly available. The compiled and formatted datasets can be provided upon valid requests.
Credit authorship contribution statement
Kanad Sen: Conceptualization, Methodology, Validation, Investigation, Coding, Writing – original draft.
Saksham Gupta: Conceptualization, Methodology, Validation, Investigation, Coding, Writing – original draft.
Abhishek Raj: Data Curation, Visualization, Validation, Investigation, Coding, Writing – original draft.
Alankar Alankar: Writing – review and editing, Validation, Supervision
Declaration of Competing Interest
The authors declare that they have no known competing financial interests or personal relationships that could have appeared to influence the work reported in this paper.